A can do a certain job in 40days which b alone can do it in 30 days .A started the work and was joined by after 19 days .In how many days was the work completed?
Answer by flame8855(424)

rate (A)= 1/40
Rate (B)= 1/30
if together = 1/40+1/30 = 7/120=7/120
for 19 days
19/40 will be done and only 21/40 will be left
7x/120=21/40
x= 9
total days = 19+9=28 days